OpenStudy (anonymous):

Solve 6 radical x+1^ – 3 = 15 A.x = 8 B. x = 107 C. x = 71 D.x = 75

OpenStudy (anonymous):

I don't get the question... can u write the equation?

OpenStudy (anonymous):

\[6\sqrt{x+1}^-3=15\]

OpenStudy (anonymous):

none from your options given... u can put in and check.

OpenStudy (anonymous):

or else shud be 8. if i m getting it. transfer 3 to RHs, it is 18. now x+1 = 9 (on squaring), x=8. so three shud nit be an exponent

OpenStudy (anonymous):

Thank you<3333